OpenStudy (anonymous):

completing the square x^2+7x+12=0

hero (hero):

Do you really have to complete the square?

OpenStudy (anonymous):

if you want

hero (hero):

It can be solved with out completing the square

hero (hero):

x^2 + 7x + 12 x^2 + 3x + 4x + 12 x(x+3)+4(x+3) (x+3)(x+4) That's without completing the square

OpenStudy (anonymous):

with completing the square

OpenStudy (anonymous):

x^2-2*(7/2)x+(7/2)^2+12=(7/2)^2 (x+7/2)^2=49/4-12 (x+7/2)^2=1/4 x+7/2=+-1/2 x=-7/2-1/2; x=-7/2+1/2 x=-4;x=-3;

hero (hero):

Completing the square is retarded